what is the ratio of 4 to 5 if the sum is 45
Answer by checkley71(8403)

4:5 WHEN SUM=45
4X+5X=45
9X=45
X=45/9
X=5
SO THE RATIOS ARE:
4*5:5*5=20:25
PROOF
20+25=45